OpenStudy (anonymous):

Help Please . For any two integer c and f, let c @ f = c2-f 2. _______ c-f What is the value of c @f when c=3 and f=4?

OpenStudy (anonymous):

answer is 7 :p

OpenStudy (anonymous):

(3)^2 - (4)^2 ----------- 3-4 =(9-16)/-1 = 7